OpenStudy (anonymous):

find the length of a focal chord making an angle theta with the axis of parabola.

OpenStudy (shubhamsrg):

sorry for the late reply..was busy somewhere else..  i'll take the general case of y^2 =4ax and hope you'll be able to generalize it.. clearly, eqn of the line is y= tan@ (x-a) it cuts the parabola y^2 = 4ax at 2 points.. solving both, we have tan^2 @ (x-a)^2 = 4ax just solve for x from here,,you'll get 2 values..thus you'll get 2 values of y.. use distance formulla for the dist between these 2 points..